b) Search, understand and discuss the with the help of sketch the difference between Undamped and Damped free vibration.
Answer #1

Before we start let us assume that a mass 'm' is vibrating on a spring of spring constant 'k' and a viscous medium of damping coefficient 'c'. As this is a free vibration, we will do not have to deal with any disturbing force. Then using Newton's second law of motion we can write

m\ddot{x}+c\dot{x}+kx=0

where x is the position from the mean point, \dot{x} is the velocity at that point and \ddot{x} is the acceleration at that point. Now let us analyze the 2 cases 1 by one

UNDAMPED FREE VIBRATION

The salient features of an undamped vibration are as follows

  • In this case, vibration is not affected by the effect of any restricting agent, like the medium or a damper.
  • If undisturbed by any external agent, the vibration tends to go on for an indefinite time.
  • As evident, such a model is impossible to be replicated in the real world. Even if we try to do it in a vacuum chamber, eventually the energy will degrade causing it to stop, though it will take a long time.

For undamped vibration, the damping constant c is 0. Thus the equation of oscillation reduces to

mi + ks = 0

or, c + (k/m).x = 0

From here, we can obtain the following important equations

  • the natural frequency of vibration w = \sqrt{k/m}~~rad/s
  • the natural frequency of vibration f = 2*\pi /w=2\pi\sqrt{k/m}~Hz
  • The time period of vibration T = 1/f =\frac{1}{2\pi}\sqrt{m/k}~s

Also solving the ODE for the equation of motion, we get

x = Asin(wt)+Bcos(wt)

where A and B are constants, which are to be determined by the initial conditions for position and velocity of vibrating particle

DAMPED FREE VIBRATION

The salient features of this type of vibration are

  • In this case, vibration dies out due to internal molecular friction of vibrating mass and friction of medium of vibration.
  • The effect o0f decay of the vibration amplitude is called "damping"
  • As evident, this is a lot more closer to the real world, and literally almost every vibration we see in our surroundings are viscous, meaning they needed a continuous supply of force to continue vibrating, in absence of which the vibration stops.

The equation given at first is the equation for damped free vibrations.

m\ddot{x}+c\dot{x}+kx=0

or,~\ddot{x}+(c/m)\dot{x}+(k/m)x=0

The roots of this equation are

x=-\frac{c}{2m}\pm \sqrt {(c/2m)^2 - (s/m)}

From here we get the damping ratio as \xi = \sqrt{\frac{(c/2m)^2}{(s/m)}} =\frac{c}{2 \sqrt{sm}}

Depending on the value of the damping ratio, there are 3 cases

  • Over damping. In this case, the damping ratio is more than 1. As evident, the damping force is stronger than the spring force, and hence the vibration is not possible.
  • Critical damping. In this case, the value of the damping constant is 1. In this case, also vibration is not possible.
  • Under damping. In this case, the value of damping constant is less than 1. vibration dies out slowly due to decay

Please note that undamped vibration is also a special case of damping in which the value of the damping ratio is 0.

The solution of the differential equation for damped vibration is given by

x = Ae^{\alpha_1t}+Be^{\alpha_2t}

where \alpha_1 and \alpha_2 are the roots of the equation, and A and B are constants determined by the initial conditions.

Plotting these motions on a graph, we obtain something like this (here free vibrations mean free undamped vibrations).

[critical damp! G= 1 [over damp & Znatet t I 4 = 0 [under damped (free vibrations

*A humble request* - If you have any doubt, please use the comment section to communicate. This will clarify your doubt, and also help me to get better at answering your next questions. At the same time, If my answer helped you, please consider leaving an upvote. I hope you understand my viewpoint. Thank you :)